OpenStudy (anonymous):

Jake is comparing the prices of two mattress cleaning companies. Company A charges $30 per mattress and an additional $13 as service charges. Company B charges $28 per mattress and an additional $15 as service charges. Part A: Write equations to represent Company A and Company B's total mattress cleaning charges for a certain number of mattresses. Define the variable used in the equations. Part B: Which company would charge less for cleaning 4 mattresses? Justify your answer. Part C: How much money is saved by using the services of Company B instead of Company A to clean 7 mattresses

OpenStudy (anonymous):

Comp A- 30m + 13= c ( c is cost) Comp B- 28m + 15= c ( m is mattresses) Company B is cheaper, just plug and chug the solution 4 into the equation! $27 dollars are saved by using Company B's service than A!

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

No prob!